Prove for any ring R and a,b∈ R , (a+b)²= a²+2ab+b²

1
Expert's answer
2021-11-18T11:01:12-0500

Any commutative ring "R" and "a,b\in R" , "(a+b)^2=a^2+2ab+b^2"

Prove:

"(a+b)^2=(a+b)(a+b)=a(a+b)+b(a+b)=\\
=a\cdot a+a\cdot b+b\cdot a+b\cdot b=\\
=a^2+a\cdot b+b\cdot a+b^2\\
(a+b)^2=(a+b)(a+b)\\
a^2+2ab+b^2=a^2+ab+ba+b^2\\
ab=ba"

That a ring "R" is commutative